OpenStudy (anonymous):

The length of a rectangle is 4 inches more than its width, and its perimeter is 34 inches. Find the length and width of the rectangle. Find the width of the rectangle

OpenStudy (midhun.madhu1987):

Let width be x. Then length = x+4 Perimeter = 2(l + w) 34 = 2 (x+ x + 4) find x.